How long will it take to fill a sphere-shaped vat with radius 27.7 ft with a certain liquid (d=2.13 g/mL) if it is spilling out of the hose at 12005.3 g/s?
(1.00 in = 2.54E0 cm)
1
Expert's answer
2019-09-24T05:57:40-0400

Solution:

Radius in meters:

"R=27.7 ft=8.44 m"

Volume of the vat:

"V=4/3 (\pi R^3)=4/3 (3.14×(8.44m)^3)=2518 m^3=2.518×10^6 L"

Mass of fluid:

"m=V×\rho=2.518×10^6 L×2130 g/L=5.363×10^9 g"

Estimated time:

"t=m\/Q=5.363\u00d710^9g \/12005.3g\/s=44.673\u00d710^4 s\u22485.17 days"